# Towards Disentangled Preference Optimization Dynamics Beyond Likelihood Displacement

### 25 Apr 2026 v2

@@ Abstract
Preference optimization is widely used to align large language models (LLMs) with human preferences. However, many margin-based objectives suppress the chosen response along with the rejected one, a phenomenon known as likelihood displacement, and no general mechanism currently prevents this across objectives. We bridge this gap by presenting a unified incentive-score decomposition of preference optimization, revealing that diverse objectives share identical local update directions and differ only in their scalar weighting coefficients. Building on this decomposition, by analyzing the dynamics of the chosen/rejected likelihoods, we identify the disentanglement band (DB), a simple, testable condition that characterizes when training can avoid likelihood displacement by realizing the preferred pathway: suppressing the loser while maintaining the winner, possibly after an initial transient. Leveraging the DB, we propose a plug-and-play reward calibration (RC) that adaptively rebalances chosen versus rejected updates to satisfy the DB and mitigate likelihood displacement, without redesigning the base objective. Empirical results show that RC steers training toward more disentangled dynamics and often improves downstream performance across a range of objectives. Our code is available at https://github.com/IceyWuu/DisentangledPreferenceOptimization.
@@
Figure 1: Overview of diagnosis and intervention for disentangled preference optimization dynamics. Left: Entangled objectives can couple winner and loser updates, leading to Pathway (i) (both likelihoods increase) or Pathway (ii) (both decrease). Upper Right: The disentanglement band (DB) provides a local, testable condition for entering Pathway (iii) (winner not down, loser not up), often after an initial transient. Lower Right: Reward calibration (RC) is a plug-and-play intervention that rescales backward gradients without changing the forward pass, making training more likely to fall into the DB and exhibit Pathway (iii), without redesigning the base objective.
@@ 1 Introduction
We address this with an incentive-score decomposition. Regardless of objective form, the gradient update decomposes along the same score directions; objective design mainly enters through two scalar incentives that weight them (Tab. 1). This yields a continuous-time analysis (Sec. 3.2) and a simple, testable condition—the disentanglement band (DB)—that characterizes when training can enter Pathway (iii) (Tab. 2), possibly after an initial transient.
@@
    We introduce an incentive-score decomposition that applies to both entangled margin-based and disentangled density ratio-based objectives, isolating scalar incentives as the key control variables (Tab. 1).
@@
    We derive likelihood dynamics and identify the DB, a simple, testable condition that characterizes when training can avoid likelihood displacement and realize the desired Pathway (iii) regime.
@@
    We propose a plug-and-play RC wrapper that rebalances chosen vs. rejected updates to mitigate likelihood displacement and promote Pathway (iii) across diverse objectives, without redesigning the base objectives.
@@
    We empirically verify that RC improves likelihood dynamics across objectives and model scales, often improving downstream alignment on benchmarks and in open-ended evaluation (Tab. 4, Sec. C.4).
@@ 3.1 Unifying Objectives via Incentives
Pathway (iii) is generally preferable (Yuan et al., 2025) because it suppresses the rejected response while preserving the chosen one, thereby mitigating the likelihood displacement problem (Razin et al., 2025).
@@ 3.2 Likelihood Dynamics and Pathway Conditions
In practice, scores are typically positively correlated ($\langle{\bm{s}}_{w,t},{\bm{s}}_{l,t}\rangle>0$, as illustrated in Fig. 3(a)). If the objective assigns too much weight to the rejected incentive, e.g., $d_{l,t}=d_{w,t}$ as in DPO, the drag can dominate, causing $\dot{z}_{w,t}<0$ even when the desired behavior is to avoid driving $z_{w,t}$ downward. This explains the squeezing effect (Ren and Sutherland, 2025). These dynamics are local-in-time, so training may enter the desired regime only after an initial transient, consistent with our empirical results (Fig. 2(a)).
@@
Figure 3: (a) Cosine similarity $\rho_{t}$ between score directions ${\bm{s}}_{w,t}$ and ${\bm{s}}_{l,t}$; larger $\rho_{t}$ indicates stronger winner-loser coupling and a narrower DB in Eq. 9. (b) Corresponding likelihood trajectories $(z_{w,t},z_{l,t})$, indicating whether training reaches the Pathway (iii) regime ($z_{w,t}$ not down, $z_{l,t}$ not up), or follows entangled Pathways (i)/(ii). (c)(d) A wide DB alone is insufficient: even with a wide DB (small $\rho_{t}$), training can remain entangled when $\log r_{t}$ stays outside the band (CPO, Fig. 3(c)). In contrast, keeping $\log r_{t}$ inside the DB is closely associated with entering the Pathway (iii) regime in our experiments (DIL-BCE, Fig. 3(d)). See Sec. C.5 for full results across models (Pythia-410M, Pythia-2B, Mistral-7B, and Qwen2.5-7B) and objectives.
@@
$\left[\rho_{t}\tfrac{\|{\bm{s}}_{l,t}\|}{\|{\bm{s}}_{w,t}\|},\tfrac{1}{\rho_{t}}\tfrac{\|{\bm{s}}_{l,t}\|}{\|{\bm{s}}_{w,t}\|}\right]$ $(\geq 0,\leq 0)$ (iii) $z_{w,t}$ not down, $z_{l,t}$ not up
@@
  • Inside DB: The incentive ratio $r_{t}$ is balanced to overcome the correlation $\rho_{t}$. The update suppresses the loser without driving the winner downward (Pathway (iii)).
@@ Conclusion.
We presented a unified framework for diagnosing and mitigating likelihood displacement in preference optimization. Through an incentive-score decomposition, we showed that diverse objectives share identical update directions and differ only in scalar incentive coefficients. This led to the disentanglement band (DB), a testable condition that characterizes when training avoids likelihood displacement by entering the Pathway (iii) regime (chosen not down, rejected not up, possibly after an initial transient). To enforce this condition, we introduced reward calibration (RC), a plug-and-play wrapper that rebalances chosen vs. rejected updates without redesigning the base objective. Empirically, RC often steers training toward Pathway (iii) and improves downstream alignment across diverse objectives and model scales.
@@ C.5 Additional Results on Preference Optimization Dynamics w/ or w/o RC
For CPO on Mistral-7B (Fig. 11(c)), RC still centers the realized log-ratio near $\log r_{t}^{\star}$ and keeps it largely within DB. In this setting, the likelihood traces can exhibit short intervals where both likelihoods move downward together, which is consistent with our regime-based interpretation: DB characterizes an instantaneous feasibility condition, and the desired Pathway (iii) behavior is best understood as a training-time regime that the trajectory can enter and revisit (often after an initial transient), rather than a guarantee of monotone per-step drift over the entire run. Empirically, with RC the trajectory spends substantially more of training in the DB-feasible region and shows a more Pathway (iii)-like profile overall, compared to the uncalibrated baseline.

### 01 May 2026 v3

# Towards Disentangled Preference Optimization Dynamics: Suppress the Loser, Preserve the Winner

@@ Abstract
Preference optimization is widely used to align large language models (LLMs) with human preferences. However, many margin-based methods also suppress the chosen response when they try to suppress the rejected one, and there is no general way to prevent this across different objectives. We address this issue with a unified incentive-score decomposition of preference optimization, revealing that different objectives share the same local update directions and differ only in their scalar weights. This decomposition provides a common framework for analyzing objectives that were previously studied in separate settings. Building on this decomposition, by analyzing the dynamics of the chosen/rejected likelihoods, we identify the disentanglement band (DB), a simple, testable condition that tells us when training can follow the desired path: suppress the loser while preserving the winner, possibly after an early stage. Using the DB, we propose reward calibration (RC), a plug‑and‑play method that adaptively rebalances the updates for chosen and rejected responses to satisfy the DB, without redesigning the base objective. Empirical results show that RC leads to more disentangled dynamics, with better downstream performance observed across several settings. Our code is available at https://github.com/IceyWuu/DisentangledPreferenceOptimization.
@@
Figure 1: Overview of diagnosis and intervention for disentangled preference optimization dynamics. Left: Entangled objectives can couple winner and loser updates, leading to Pathway (i) (both likelihoods increase) or Pathway (ii) (both decrease). Upper Right: The disentanglement band (DB) provides a local, testable condition for entering Pathway (iii) (suppress the loser, preserve the winner). Lower Right: Reward calibration (RC) is a plug-and-play intervention that rescales backward gradients without changing the forward pass, making training more likely to fall into the DB and exhibit Pathway (iii), without redesigning the base objective.
@@ 1 Introduction
We address this with an incentive-score decomposition. Regardless of objective form, the gradient update decomposes along the same score directions; objective design mainly enters through two scalar incentives that weight them (Tab. 1). This yields a continuous-time analysis (Sec. 3.2) and a simple, testable condition—the disentanglement band (DB)—that characterizes when training can enter Pathway (iii) (Tab. 2), possibly after an early stage.
@@
    We introduce an incentive-score decomposition that unifies entangled and disentangled objectives, isolating scalar incentives as key controls (Tab. 1).
@@
    We derive likelihood dynamics and identify the DB, a simple condition for the desired Pathway (iii): suppress the loser while preserve the winner.
@@
    We propose plug‑and‑play RC, which rebalances chosen vs. rejected updates to promote Pathway (iii) without redesigning the base objective.
@@
    We verify that RC improves likelihood dynamics across objectives and scales, with downstream improvements in several settings (Tab. 4, Sec. C.4).
@@ 3.1 Unifying Objectives via Incentives
Pathway (iii) is a natural desideratum in the post-SFT regime (Yuan et al., 2025), as it avoids likelihood displacement and gradient entanglement without degrading the chosen response. Although not universally optimal, it provides a principled direction for training dynamics.
@@ 3.2 Likelihood Dynamics and Pathway Conditions
In practice, scores are typically positively correlated ($\langle{\bm{s}}_{w,t},{\bm{s}}_{l,t}\rangle>0$, as illustrated in Fig. 3(a)). If the objective assigns too much weight to the rejected incentive, e.g., $d_{l,t}=d_{w,t}$ as in DPO, the drag can dominate, causing $\dot{z}_{w,t}<0$ even when the desired behavior is to avoid driving $z_{w,t}$ downward. This explains the squeezing effect (Ren and Sutherland, 2025). These dynamics are local-in-time, so training may enter the desired regime only after an early stage, consistent with our empirical results (Fig. 2(a)).
@@
Figure 3: (a) Cosine similarity $\rho_{t}$ between score directions ${\bm{s}}_{w,t}$ and ${\bm{s}}_{l,t}$; larger $\rho_{t}$ indicates stronger winner-loser coupling and a narrower DB in Eq. 9. (b) Corresponding likelihood trajectories $(z_{w,t},z_{l,t})$, indicating whether training reaches the Pathway (iii) regime (suppress $z_{l,t}$, preserve $z_{w,t}$), or follows entangled Pathways (i)/(ii). (c)(d) A wide DB alone is insufficient: even with a wide DB (small $\rho_{t}$), training can remain entangled when $\log r_{t}$ stays outside the band (CPO, Fig. 3(c)). In contrast, keeping $\log r_{t}$ inside the DB is closely associated with entering the Pathway (iii) regime in our experiments (DIL-BCE, Fig. 3(d)). See Sec. C.5 for full results across models (Pythia-410M, Pythia-2B, Mistral-7B, and Qwen2.5-7B) and objectives.
@@
$\left[\rho_{t}\tfrac{\|{\bm{s}}_{l,t}\|}{\|{\bm{s}}_{w,t}\|},\tfrac{1}{\rho_{t}}\tfrac{\|{\bm{s}}_{l,t}\|}{\|{\bm{s}}_{w,t}\|}\right]$ $(\geq 0,\leq 0)$ (iii) suppress $z_{l,t}$, preserve $z_{w,t}$
@@
  • Inside DB: The incentive ratio $r_{t}$ is balanced to overcome the correlation $\rho_{t}$. The update suppresses the loser and preserves the winner (Pathway (iii)).
@@ Conclusion.
We presented a unified framework for diagnosing and steering the reward dynamics of preference optimization. Through an incentive-score decomposition, we showed that diverse objectives share identical update directions and differ only in scalar incentive coefficients. This decomposition itself provides a common analytical lens for comparing objectives that have previously been studied in separate frameworks. This led to the disentanglement band (DB), a testable condition that characterizes when training enters the Pathway (iii) regime: suppress the loser while preserving the winner, possibly after an early stage. To enforce this condition, we introduced reward calibration (RC), a plug-and-play wrapper that rebalances chosen vs. rejected updates without redesigning the base objective. Empirically, RC often steers training toward Pathway (iii), with downstream improvements observed in several settings across diverse objectives and model scales.
@@ C.5 Additional Results on Preference Optimization Dynamics w/ or w/o RC
For CPO on Mistral-7B (Fig. 11(c)), RC still centers the realized log-ratio near $\log r_{t}^{\star}$ and keeps it largely within DB. In this setting, the likelihood traces can exhibit short intervals where both likelihoods move downward together, which is consistent with our regime-based interpretation: DB characterizes an instantaneous feasibility condition, and the desired Pathway (iii) behavior is best understood as a training-time regime that the trajectory can enter and revisit (often after an early stage), rather than a guarantee of monotone per-step drift over the entire run. Empirically, with RC the trajectory spends substantially more of training in the DB-feasible region and shows a more Pathway (iii)-like profile overall, compared to the uncalibrated baseline.